OpenStudy (anonymous):

Find the missing value to the nearest hundreth. Sin______=9/16 Not sure how to find this DX Any explanations would be greatly appreciated!!

OpenStudy (anonymous):

Let A be the unknown. sin A = 9/16 when we convert 9/16 to decimal sin A = 0.5625 \[A = \frac{ 0.5625 }{ \sin }\] which is equal to \[A = \sin^{-1} (0.5625)\] Using your calculator, A = 34.23

OpenStudy (anonymous):

Ok that makes sense